Exeter has a material standard of 1 pound per unit of output. Each pound has a standard price of $27 per pound. During July, Exeter paid $141,000 for 4,970 pounds, which they used to produce 4,760 units. What is the direct materials quantity variance?

$5,910 favorable

$6,810 unfavorable

$8,270 unfavorable

$5,670 unfavorable
Answer #1

Materials Quantity Variance = (SQ-AQ)*SP

= (4760*1 - 4970) * 27

= (4760 - 4970) * 27

= 5670 Unfavorable

Option D is the answer